OpenStudy (anonymous):

Sam has some quarters, nickels, and dimes. He has 4 more quarters than dimes and 3 more dimes than nickels. If n represents the number of nickels he has, which of the following represents, in cents, the total value of all his coins? A) 40n+205 B) 40n+130 C) 40n+ 7 D) 7n+130 E) 3n+10

OpenStudy (anonymous):

wow... this one is "fun"

OpenStudy (anonymous):

if n = nickels and he has 3 more dimes than nickels, then he has (n + 3) dimes, and if he has 4 more quarters than dimes, then he has (n + 3 + 4) quarters, or (n + 7) quarters. nickels = n dimes = n + 3 quarters = n + 7 since nickels are worth 5 cents, multiply n x 5 since dimes are worth 10 cents, multiply (n + 3) x 10, and since quarters are worth 25 cents, multiply (n + 7) x 25 so 5n + 10(n + 3) + 25(n + 7) will give you your final answer.